In flies, body (e) is recessive to the WT (E). In the mating CcDdEe x ccddee You observe the following offspring: curled wings (c) is recessive to the WT (C. Cardinal eyes (d) is recessive to the WT (D). Ebony 88 wild type wings, wild type bodies, wild type eyes doubta Cc. ddee 4 cardinal eyes singt Ce bd ee 658 ebony bodies cc dd ee 50 ebony bodies and cardinal eyes Sinao cc pd Ee52 curled wings Sing ccdd Ee 612 curled wing and cardinal eyes e Dd ee 10 curled wings and ebony bodies Poubie e singta dl e cde 84 curled wings, ebony bodies, and cardinal eyes Ccddee c ddee c cc bd Ee ce ddEe 8) which are the parental classes? CcDdee. Single crossovers? Pond codies ebnCardinal , curle d u tedcardns c do Ee cc Ddee epon Double crossovers? 9) What is the gene order? 10) Draw a map of the genes including the distances between them
Answer #1

Answer:

8).

Parental classes:

Ebony bodies

Curled wings and cardinal eyes

Single crossovers:

Wild-type (all)

Curled wings, ebony bodies, and cardinal eyes

Ebony bodies, and cardinal eyes

Curled wings

Double crossovers:

Cardinal eyes

Curled wings, and ebony bodies.

9). e-c-d

10). e--------11.94mu------c----7.45mu----d

Explanation:

In flies, body (e) is recessive to the WT (E). In the mating CcDdEe x ccddee You observe the following offspring: curled wings (c) is recessive to the WT (C. Cardinal eyes (d) is recessive to the WT (D). Ebony 88 wild type wings, wild type bodies, wild type eyes 4 cardinal eyes 658 ebony bodies 50 ebony bodies and cardinal eyes 52 curled wings - 612 curled wing and cardinal eyes - 10 curled wings and ebony bodies 84 curled wings, ebony bodies, and cardinal eyes CED CEd CeD Ce d cED c E d ce D c e d Ccddee c ddee 8) Which are the parental classes? C cc bd Ee ce ddEe CcDdee. Single crossovers? Pond codies ebnCardinal , curle d u tedcardns c do Ee cc Ddee Double crossovers? 9) What is the gene order? 10) Draw a map of the genes including the distances between them

Hint: Non-recombinant (parental) combinations are produced more in number than recombinants. The parental (non-recombinant) genotypes is c E d / C e D

1).

If single crossover occurs between c&e

Normal combination: cE/Ce

After crossover: ce/CE

ce progeny=10+84 = 94

CE progeny = 4+88 = 92

Total this progeny = 186

Total progeny = 1558

The recombination frequency between c&e = (number of recombinants/Total progeny) 100

RF = (186/1558)100 = 11.94%

2).

If single crossover occurs between e&d

Normal combination: eD/Ed

After crossover: cv+ ed/ED

ed progeny= 50+84 = 134

ED progeny = 52+88 = 140

Total this progeny = 274

The recombination frequency between e&d = (number of recombinants/Total progeny) 100

RF = (274/1558)100 = 17.59%

3).

If single crossover occurs between c&d

Normal combination: CD/cd

After crossover: Cd/cD

Cd progeny=50+4=54

cD progeny = 52+10=62

Total this progeny = 116

Total progeny = 1558

The recombination frequency between c&d = (number of recombinants/Total progeny) 100

RF = (116/1558)100 = 7.45%

Recombination frequency (%) = Distance between the genes (mu)

e--------11.94mu------c----7.45mu----d